OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 14 when x = 8. A. 112 B. 2 C. 7/4 D. 7

OpenStudy (mertsj):

y=k/x

OpenStudy (mertsj):

Plug in the given y and x values and find k which is the constant of variation.

OpenStudy (anonymous):

I don't get it

OpenStudy (mertsj):

Did you see the equation I wrote?

OpenStudy (anonymous):

yes I did so it would be 14=k/8

OpenStudy (mertsj):

\[y=\frac{k}{x}\]

OpenStudy (mertsj):

Now solve that for k by multiplying both sides by 8

OpenStudy (mertsj):

\[14(8)=\frac{k}{8}(8)\]

OpenStudy (anonymous):

112=k/64?

OpenStudy (mertsj):